### v2

# Regularity of solutions to degenerate and singular free boundary problems with volume constraint

Abstract.
We prove existence and regularity of solutions to degenerate and singular elliptic free boundary problems, where the volume of the positivity set of the solution is prescribed.
Key words and phrases:
Volume constrained free boundary problems, degenerate and singular elliptic PDEs, regularity of solutions
2010 Mathematics Subject Classification:
35R35, 35B65, 42B37
The third author was partially supported by Simons Foundation grant MP-TSM-00002709
1. Introduction
Our aim is to show existence and regularity of solutions $u$ to the following weighted volume constrained free boundary problem, where the measure of the positivity set of $u$ is prescribed. Let $\omega=\omega(x)$ be a Muckenhoupt $A_{2}$ weight on $\mathbb{R}^{n}$, $n\geq 2$. For a measurable set $E$ we denote $\omega(E)=\int_{E}\omega(x)\,dx$. Fix $0<m<\omega(B_{1})$ and a smooth function $g$ on $\overline{B_{1}}$ such that $g\geq c>0$ on $\partial B_{1}$. We consider the weighted minimization problem
(1.1) $\min_{v\in K_{0}}\int_{B_{1}}|\nabla v|^{2}\omega\,dx$
where
$K_{0}=\big\{v\in H^{1}(B_{1},\omega):\,v\big|_{\partial B_{1}}=g,\,\omega(\{v>0\}\cap B_{1})=m\big\}$
and $H^{1}(B_{1},\omega)$ denotes the corresponding weighted Sobolev space (see Section 2 for notation).
Free boundary problems with volume constraints as (1.1) have received significant attention over the past five decades owing to their crucial role in applied sciences and the mathematical challenges they present (see, for example, [AguiAltCaff86, ACS87, BMW06, Teix05, Teix10]). These problems arise in the study of best insulation devices, semiconductor theory and plasma physics, just to name a few. From a mathematical perspective, the questions of existence of an optimal configuration $u$, as well as regularity properties of $u$ and of the free boundary $\partial\{u>0\}$, pose important difficulties. The first, most influential work in this direction is the paper by Aguilera–Alt–Caffarelli [AguiAltCaff86], in which they consider (1.1) with no weight, that is, $\omega\equiv 1$. In this case, they reduce the problem to a one phase free boundary problem and apply the results of Alt–Caffarelli [ALtCaff81] to prove existence and regularity of solutions. Generalizations of the unweighted case were carried out in, for example, [Led96, AFMT99, Tilli99]. More recently, nonlocal versions of the unweighted problem were considered in, for instance, [TeixTeym, DaSilvaRossi19]. In particular, in [TeixTeym], the Caffarelli–Silvestre extension problem for the fractional Laplacian in $\mathbb{R}^{n}$ (see, for instance, [Stinga_2024, Part IV]) leads the authors to consider weighted equations of the form $\mathrm{div}(\omega(X)\nabla u)=f(X)$, but the problem is different than (1.1). Our work is thus motivated by those of Aguilera–Alt–Caffarelli [AguiAltCaff86] and Teixeira–Teymurazyan [TeixTeym].
From the mathematical viewpoint, this article is the first (up to the best of our knowledge) to establish the existence of an open optimal set $\{u>0\}$ with prescribed volume in weighted spaces, and the regularity of $u$. Our problem (1.1) presents at least two challenges compared to other optimal design problems governed by elliptic equations. The first one concerns the local regularity of minimizers. Since $A_{2}$ weights are, in general, only measurable, locally integrable functions, one can only expect Hölder continuity for a minimizer $u$. This limitation prevents us from applying the ideas and techniques used in [AguiAltCaff86], [ALtCaff81] or [TeixTeym] to prove existence and regularity, where the Lipschitz regularity of minimizers up to the free boundary $\partial\{u>0\}$ plays a fundamental role. The second difficulty is the absence of an isoperimetric inequality for general $A_{2}$ weights, which is an essential tool in our approach for adjusting the measure of the positivity set to the desired value in order to prove existence. To be more precise, in the proof of existence, we need to assume that $\omega\in A_{2}$ satisfies either one of the following two conditions.
  • Weighted isoperimetric inequality. There is a constant $C_{0}>0$ such that, for any $u\in W^{1,1}(B_{1})$ and $s>0$,
    (1.2) $\omega(\{u>s\}\cap B_{1})^{\frac{n-1}{n}}\leq C_{0}\int_{\{u=s\}\cap B_{1}}\omega\,d\mathcal{H}^{n-1}$
    where $d\mathcal{H}^{n-1}$ denotes the (n1)(n-1)-dimensional Hausdorff measure in $\mathbb{R}^{n}$.
  • Bound from below. There is $\tau>0$ such that
    (1.3) $\omega(x)\geq\tau\qquad\hbox{for a.e.}~x\in B_{1}.$
However, only the $A_{2}$ condition on $\omega$ is used to prove regularity.
Our main result reads as follows (see Section 2 for notation).
Theorem 1.1 (Existence and regularity of solutions).
Let $\omega$ be an $A_{2}$ weight. Assume that either $\omega$ satisfies the isoperimetric inequality (1.2), or that it is bounded below away from zero in $B_{1}$ as in (1.3). Then there is a minimizer $u\in K_{0}$ to problem (1.1) such that $0\leq u\leq\|g\|_{L^{\infty}(\partial B_{1})}$ in $B_{1}$. Moreover, $u$ is locally Hölder continuous in $B_{1}$, that is, there exists $0<\beta<1$ depending only on $\omega]_{A_{2}}$ and $n$ such that, for any compact set $K\subset B_{1}$,
$u]_{C^{\beta}(K)}\leq C\|u\|_{L^{2}(B_{1},\omega)}$
where $C>0$ depends only on $K$, $\omega]_{A_{2}}$ and $n$. Furthermore, $u$ is a weak solution to $\mathrm{div}(\omega(x)\nabla u)=0$ in the open set $\{u>0\}\cap B_{1}$.
It is an open problem to characterize the weights $\omega$ for which (1.2) is true and, in particular, whether or not is true for general $A_{2}$ weights.
Examples of $A_{2}$ weights for which (1.2) holds are $A_{1}$ weights, that is, weights $\omega$ for which there is a constant $C>0$ such that $M\omega(x)\leq C\omega(x)$ for a.e. $x\in\mathbb{R}^{n}$, where $M$ is the classical Hardy–Littlewood maximal function (note that $A_{1}\subset A_{2}$). Indeed, this is a consequence of the results in [Turesson00, Section 2.4.2]. In fact, (1.2) is stated as [Turesson00, Theorem 2.4.8] (this result is valid for our case because Theorem 2.4.7–the co-area formula, and Lemma 2.4.1–the Poincaré inequality, in [Turesson00] are true under our hypotheses as well). Furthermore, we do not require the $A_{1}$ weight to be continuous. Canonical examples of $A_{2}$ weights satisfying (1.3) are $\omega(x)=|x|^{-\alpha}$ for $0<\alpha<n$, or, more generally, $\omega(x)=\operatorname{dist}(x,Z)^{-\alpha}$, where $Z$ is a “thin” set of zero Lebesgue measure, and anisotropic variations of power weights (see [LST] where the lower bound in $\omega$ is assumed to study a weighted one phase problem generalizing [ALtCaff81]). In particular, for the case of the extension problem for the fractional Laplacian $(-\Delta)^{s}$, where $\omega(x,y)=|y|^{1-2s}$ for $(x,y)\in\mathbb{R}^{n+1}$, we have the restriction $1/2\leq s<1$.
Our approach to the proof of Theorem 1.1 relies on a penalization method that allows us to study the problem without relying on the smoothness of the free boundary as in [AguiAltCaff86]. The new functional to be minimized takes the form
$\int_{B_{1}}|\nabla v|^{2}\omega\,dx+f_{\varepsilon}(v)$
for some parameter $\varepsilon>0$. Assuming only the $A_{2}$ condition on the weight, we show that minimizers of the penalized problem exist for all $\varepsilon>0$ and are Hölder continuous. The Hölder continuity is proved using a blow-up approximation by constants in the spirit of Caffarelli’s regularity technique. Notice that this technique was also implemented in [LST] for the one-phase problem. However, unlike [LST], we are able to establish Hölder continuity of solutions to our problem (1.1) without imposing any further restrictions on the weight $\omega$ (other than being in the $A_{2}$ class). Then we show that there exists $\varepsilon_{\ast}>0$ small enough such that $u_{\ast}$ is a minimizer for the penalized problem if and only if $u_{\ast}$ is a minimizer for (1.1). Here we argue by contradiction, finding an appropriate competitor function $u_{t}$ for which $\{u_{t}>0\}$ approximates in measure $\{u>0\}$, and we apply the hypothesis on $\omega$ stated in Theorem 1.1. This argument for proving existence of minimizers was implemented in [BHP05] in the unweighted case, where Lipschitz continuity of minimizers was also proved.
The analysis presented in this paper extends without difficulty to volume constrained free boundary problems involving more general energies of the form
$\int_{\Omega}(\nabla v)^{T}A(x)\nabla v\,dx$
where $\Omega$ is a smooth bounded domain and the matrix of coefficients $A(x)$ is bounded, measurable, and degenerate elliptic with respect to the weight $\omega$, that is, there are constants $0<\lambda\leq\Lambda$ such that $\lambda\omega(x)|\xi|^{2}\leq\xi^{T}A(x)\xi\leq\Lambda\omega(x)|\xi|^{2}$ for a.e. $x\in\Omega$ and all $\xi\in\mathbb{R}^{n}$. In this case, the constants in Theorem 1.1 will also depend on $\lambda$ and $\Lambda$. This general case is particularly interesting because it arises when considering fractional elliptic equations in divergence form $(-\mathrm{div}(a(x)\nabla))^{s}w=f$ in $\Omega$, see [Caffarelli-Stinga, Stinga-Zhang].
An interesting question is whether existence of minimizers fails if one removes either assumption (1.2) or (1.3). We believe that one should be able to remove them, but the technique will have to be different.
The paper is organized as follows. In Section 2, we introduce preliminary results about weighted Sobolev spaces and degenerate/singular elliptic equations. Section 3 is dedicated to the analysis of the penalized problem, including existence, boundedness and local regularity. These properties are derived without any additional assumptions on the $A_{2}$ weight. As mentioned earlier, to adjust the measure of the positivity set to the prescribed value and close the existence part of Theorem 1.1, we must consider weights for which either an isoperimetric inequality holds or that are bounded below away from zero in $B_{1}$. Thus, in Section 4, under such assumptions, we finalize the proof of our main result.
2. Preliminaries
In this section we collect results regarding weighted Sobolev spaces and degenerate elliptic equations. References here are [FKS, HKM, Turesson00], see also [FJK, FKJ].
We denote by $B_{R}(x)$ the ball in $\mathbb{R}^{n}$ of radius $R$ centered at $x$. When the center is the origin, we omit it and just write $B_{R}$.
A nonnegative measurable function $\omega=\omega(x)$ defined on $\mathbb{R}^{n}$ is an $A_{2}$ weight if both $\omega$ and $\omega^{-1}$ are locally integrable and
(2.1) $\omega]_{A_{2}}:=\sup\limits_{B\subset\mathbb{R}^{n}}\left(\frac{1}{|B|}\int_{B}\omega\,dx\right)\left(\frac{1}{|B|}\int_{B}\omega^{-1}\,dx\right)<\infty$
where the supremum is taken over all balls $B\subset\mathbb{R}^{n}$. Here $|E|$ denotes the Lebesgue measure of a measurable set $E\subset\mathbb{R}^{n}$. In this case, we call (2.1) the $A_{2}$ condition and write $\omega\in A_{2}$.
Throughout this section, we assume that $\omega\in A_{2}$.
By definition of $A_{2}$ weights, we may assume that $0<\omega<\infty$ almost everywhere. Moreover, as $\omega\in L^{1}_{\mathrm{loc}}(\mathbb{R}^{n})$, the induced measure defined by
$\omega(E):=\int_{E}\omega(x)\,dx$
is a positive Radon measure. Moreover, the weight $\tilde{\omega}(x):=\omega(x_{0}+rx)$, for fixed $x_{0}\in\mathbb{R}^{n}$ and $r>0$, is also an $A_{2}$ weight with $\tilde{\omega}]_{A_{2}}=[\omega]_{A_{2}}$.
Lemma 2.1 (Properties of $A_{2}$ weights).
Let $\omega$ be an $A_{2}$ weight on $\mathbb{R}^{n}$.
  1. $(1)$
    There exists a constant $c_{1}>0$ depending only on $\omega]_{A_{2}}$ such that
    $c_{1}\left(\frac{|E|}{|B|}\right)^{2}\leq\frac{\omega(E)}{\omega(B)}$
    whenever $B\subset\mathbb{R}^{n}$ is a ball and $E\subset B$ is a measurable set.
  2. $(2)$
    There exists a doubling constant $D=D(n,[\omega]_{A_{2}})>0$ such that
    $0<\omega(B_{2r}(x_{0}))\leq D\omega(B_{r}(x_{0}))$
    for all balls $B_{r}(x_{0})\subset\mathbb{R}^{n}$.
  3. $(3)$
    There exist constants $C_{2}>1$ and $0<S_{D}\leq 1$ depending only on $n$ and $D$ such that
    $\frac{\omega(E)}{\omega(B)}\leq C_{2}\left(\frac{|E|}{|B|}\right)^{S_{D}}$
    whenever $B\subset\mathbb{R}^{n}$ is a ball and $E\subset B$ is a measurable set.
By Lemma 2.1(1)(1), we get that $\omega(E)=0$ if and only if $|E|=0$.
The set $L^{2}(B_{1},\omega)$ is the Hilbert space of all measurable functions $u$ on $B_{1}$ such that
$\|u\|_{L^{2}(B_{1},\omega)}^{2}=\int_{B_{1}}|u|^{2}\omega\,dx<\infty.$
By Hölder’s inequality, $L^{2}(B_{1},\omega)\subset L^{1}(B_{1})$. The weighted Sobolev space $H^{1}(B_{1},\omega)$ is the Hilbert space of functions $u\in L^{2}(B_{1},\omega)$ such that the weak gradient $\nabla u$ is in $L^{2}(B_{1},\omega)$ under the obvious norm. We have that $C^{\infty}(\overline{B_{1}})$ is dense in $H^{1}(B_{1},\omega)$. Using standard arguments as in the unweighted case, one can show that the embedding $H^{1}(B_{1},\omega)\subset\subset L^{2}(B_{1},\omega)$ is compact. The space $H^{1}_{0}(B_{1},\omega)$ is the closure of $C^{\infty}_{c}(B_{1})$ with respect to the norm of $H^{1}(B_{1},\omega)$. If we fix $g\in C^{\infty}(\overline{B_{1}})$, then $H_{g}^{1}(B_{1},\omega)$ denotes the set of all $u\in H^{1}(B_{1},\omega)$ for which $u-g\in H^{1}_{0}(B_{1},\omega)$.
It follows by density and Hölder’s inequality that $H^{1}(B_{1},\omega)\subset W^{1,1}(B_{1})$, where $W^{1,1}(B_{1})$ is the classical Sobolev space of functions $u\in L^{1}(B_{1})$ with $\nabla u\in L^{1}(B_{1})$. In fact,
(2.2) $\|u\|_{W^{1,1}(B_{1})}\leq C\|u\|_{H^{1}(B_{1},\omega)}$
for some $C>0$ depending only on $n$ and $\omega]_{A_{2}}$. Similarly $H^{1}_{0}(B_{1},\omega)\subset W^{1,1}_{0}(B_{1})$. In particular, if $u\in H^{1}_{g}(B_{1},\omega)$ then $u=g$ on $\partial\Omega$ in the sense of traces in $W^{1,1}(B_{1})$.
The space $H^{1}(B_{R},\omega)$ is defined in a similar way by replacing $B_{1}$ by $B_{R}$, $R>0$. The next result is a direct consequence of [FKS, Theorem 1.2], see also [HKM, Chapter 15].
Lemma 2.2 (Weighted Poincaré inequality).
There is a constant $C_{P,\omega}>0$ depending only on $n$ and $\omega]_{A_{2}}$ such that, for any $u\in H_{0}^{1}(B_{R},\omega)$,
(2.3) $\int_{B_{R}}|u|^{2}\omega\,dx\leq C_{P,\omega}R^{2}\int_{B_{R}}|\nabla u|^{2}\omega\,dx.$
We say that $h\in H^{1}(B_{1},\omega)$ is a weak solution to the degenerate/singular elliptic equation
(2.4) $\mathrm{div}(\omega(x)\nabla h)=0\qquad\hbox{in}~B_{1}$
if
$\int_{B_{1}}\omega(x)\nabla h\nabla\varphi\,dx=0$
for all $\varphi\in H^{1}_{0}(B_{1},\omega)$. By [FKS], given $g\in H^{1}(B_{1},\omega)$, there exists a unique weak solution $h\in H^{1}(B_{1},\omega)$ to (2.4) such that $h-g\in H^{1}_{0}(B_{1},\omega)$. We have the following local boundedness and Hölder regularity estimate that was proved in [FKS].
Theorem 2.3 (Harnack inequality and Hölder regularity).
Let $h\in H^{1}(B_{1},\omega)$ be a weak solution to (2.4). If $h\geq 0$ in $B_{1}$, then $h$ satisfies an interior Harnack inequality, that is, there exists $C_{H}>0$ depending only on $n$ and $\omega]_{A_{2}}$ such that
$\sup_{B_{1/2}}h\leq C_{H}\inf_{B_{1/2}}h.$
As a consequence, $h$ is locally Hölder continuous in $B_{1}$, that is, there exist constants $\bar{C}>0$ and $0<\alpha<1$ depending only on $n$ and $\omega]_{A_{2}}$ such that
$\|h\|_{C^{0,\alpha}(B_{1/2})}\leq\bar{C}\left(\frac{1}{\omega(B_{1})}\int_{B_{1}}|h|^{2}\omega\,dx\right)^{1/2}.$
3. The penalized problem with parameter $\varepsilon>0$
In this section we prove existence, boundedness and local Hölder regularity of solutions to a penalized formulation with parameter $\varepsilon>0$. Throughout this section we fix a smooth function $g$ on $\overline{B_{1}}$ such that $g\geq c>0$ on $\partial B_{1}$, a parameter $\varepsilon>0$, a weight $\omega\in A_{2}$ and a constant $0<m<\omega(B_{1})$.
Let
$f_{\varepsilon}(t):=\frac{1}{\varepsilon}(t-m)^{+}\qquad\hbox{for}~t\geq 0$
and consider the functional
(3.1) $J_{\varepsilon}(v,\omega,B_{1}):=\int_{B_{1}}|\nabla v|^{2}\omega\,dx+f_{\varepsilon}(\omega(\{v>0\}\cap B_{1})).$
The penalized problem we consider is
($P_{\varepsilon}$) $\min\big\{J_{\varepsilon}(v,\omega,B_{1}):v\in H_{g}^{1}(B_{1},\omega)\big\}.$
Clearly, if $v\in H^{1}_{g}(B_{1})$ then $v^{+}=\max(v,0)\in H^{1}_{g}(B_{1})$ and $J_{\varepsilon}(v^{+},\omega,B_{1})\leq J_{\varepsilon}(v,\omega,B_{1})$.
Lemma 3.1 (Existence of minimizers for ($P_{\varepsilon}$)).
There is a minimizer $u\geq 0$ to problem ($P_{\varepsilon}$).
Proof.
Given any ball $B_{r}\subset B_{1}$, let $v_{r}\in H^{1}_{g}(B_{1}\setminus B_{r},\omega)$ be the unique weak solution to
(3.2) $\begin{cases}\mathrm{div}(\omega(x)\nabla v)=0&\hbox{in}~B_{1}\\ v=g&\hbox{on}~\partial B_{1}\\ v=0&\hbox{on}~\partial B_{r}.\end{cases}$
By the maximum principle, $\{v_{r}>0\}\cap B_{1}=B_{1}\setminus\overline{B_{r}}$. We extend $v_{r}$ as 0 inside of $B_{r}$, so that $v_{r}\in H^{1}_{g}(B_{1},\omega)$. Then, by choosing $r>0$ sufficiently large,
$\omega(\{v_{r}>0\}\cap B_{1})=\omega(B_{1})-\omega(B_{r})\leq m.$
Therefore,
(3.3) $0\leq J_{\varepsilon}(v_{r},\omega,B_{1})\leq C_{0}$
for some constant $C_{0}>0$ independent of $\varepsilon$ and
$0\leq\mu:=\inf\big\{J_{\varepsilon}(v,\omega,B_{1}):v\in H_{g}^{1}(B_{1},\omega)\big\}\leq C_{0}.$
Let $\{v_{j}\}_{j=1}^{\infty}\subset H_{g}^{1}(B_{1},\omega)$ be a minimizing sequence for $J_{\varepsilon}$. Since $\mu\leq J_{\varepsilon}(v_{j}^{+},\omega,B_{1})\leq J_{\varepsilon}(v_{j},\omega,B_{1})$, we can assume that $v_{j}\geq 0$. For all $j$ sufficiently large, by (3.3),
$\int_{B_{1}}|\nabla v_{j}|^{2}\omega\,dx\leq C_{0}.$
Moreover, by the weighted Poincaré inequality in Lemma 2.2,
$\int_{B_{1}}|v_{j}|^{2}\omega\,dx\leq 2\int_{B_{1}}|v_{j}-g|^{2}\omega\,dx+2\int_{B_{1}}|g|^{2}\omega\,dx\leq C(C_{0}+1)$
where $C>0$ is independent of $j$. Thus, there exists $u\in H^{1}(B_{1},\omega)$ such that, up to a subsequence, $v_{j}\to u$ weakly in $H^{1}(B_{1},\omega)$. By the compact embedding $H^{1}(B_{1},\omega)\subset\subset L^{2}(B_{1},\omega)$ and (2.2), up to a further subsequence, $v_{j}\to u$ strongly in $L^{2}(B_{1},\omega)$ and a.e. in $B_{1}$. In particular, $u\geq 0$ a.e.. Moreover, since $v_{j}-g$ weakly converges in $H^{1}_{0}(B_{1},\omega)$, and $H^{1}_{0}(B_{1},\omega)$ is convex and strongly closed, it follows that $u-g\in H^{1}_{0}(B_{1},\omega)$, so that $u\in H^{1}_{g}(B_{1},\omega)$.
Next, since $v_{j}$ converges in $\omega$-measure to $u\geq 0$ in $B_{1}$, it follows that
$\omega(\{u\neq 0\}\cap B_{1})=\omega(\{u>0\}\cap B_{1})\leq\liminf_{k\to\infty}\omega(\{v_{j}>0\}\cap B_{1}).$
Therefore, by the weak convergence of $\nabla v_{j}$ to $\nabla u$ and the fact that $f_{\varepsilon}$ is a continuous, nondecreasing function,
$J_{\varepsilon}(u,\omega,B_{1})\leq\liminf\limits_{j\to\infty}J_{\varepsilon}(v_{j},\omega,B_{1})=\mu$
and $u$ attains the minimum in ($P_{\varepsilon}$). ∎
Lemma 3.2 (Boundedness of minimizers).
If $u$ is the minimizer from Lemma 3.1 then
$\mathrm{div}(\omega(x)\nabla u)\geq 0\qquad\hbox{in the weak sense in}~B_{1}.$
In particular, by the maximum principle,
$0\leq u\leq\|g\|_{L^{\infty}(\partial B_{1})}\qquad\hbox{a.e.\,in}~B_{1}.$
Proof.
For a nonnegative function $\phi\in C_{c}^{\infty}(B_{1})$ and any $t>0$, we have $u-t\phi\in H^{1}_{g}(B_{1})$, so that, by the minimality of $u$,
$0$ $\leq\frac{1}{2t}(J_{\varepsilon}(u-t\phi,\omega,B_{1})-J_{\varepsilon}(u,\omega,B_{1}))$
$=-\int_{B_{1}}\omega\nabla u\cdot\nabla\phi\,dx+\frac{t}{2}\int_{B_{1}}|\nabla\phi|^{2}\omega\,dx$
$\quad+\frac{1}{2t}\big[f_{\varepsilon}(\omega(\{u-t\phi>0\}\cap B_{1}))-f_{\varepsilon}(\omega(\{u>0\}\cap B_{1}))\big].$
The last term above is nonpositive because $f_{\varepsilon}$ is nondecreasing. Thus, by letting $t\to 0$,
$\int_{B_{1}}\omega\nabla u\cdot\nabla\phi\,dx\leq 0$
and $u$ is a subsolution in $B_{1}$ in the weak sense. ∎
We next prove that minimizers for ($P_{\varepsilon}$) are Hölder continuous by utilizing the scaling features of the problem and showing that we can normalize the minimizer by considering a modified functional $\tilde{J}_{\varepsilon}$. First, we localize with the following definition.
Definition 3.3.
Given a ball $B$, we say that $u$ is a local minimizer in $B$ of a functional of the form
$J(v,\omega,B):=\int_{B}|\nabla v|^{2}\omega\,dx+f(\omega(\{v>0\}\cap B))$
if for every subdomain $E\subset B$, $u$ minimizes
$\int_{E}|\nabla v|^{2}\omega\,dx+f(\omega(\{v>0\}\cap E)+\omega(\{u>0\}\cap(B\setminus E)))$
among all functions $v$ for which $v-u\in H^{1}_{0}(E,\omega)$.
Note that the term $\omega(\{u>0\}\cap(B\setminus E))$ is a constant in the minimization in Definition 3.3 and that a minimizer $u$ of $J_{\varepsilon}(u,\omega,B_{1})$ is a local minimizer in $B_{1}$. Indeed, if there were a set $E$ and a function $v$ on $E$ for which the above functional is smaller than for $u$, then one could take $w=u$ in $B\setminus E$ and $w=v$ in $E$ to get $J_{\varepsilon}(w,\omega,B_{1})<J_{\varepsilon}(u,\omega,B_{1})$. By taking $E=B_{1}$, any local minimizer is a (global) minimizer, however it may not satisfy our imposed boundary condition so we do not use this latter property.
For $x_{0}\in B_{1/2}$, $0<r<1/2$ and $\delta>0$, we define the rescaled function
$\tilde{u}(y)=\frac{u(x_{0}+ry)}{\delta}\qquad\hbox{for}~y\in B_{1}.$
Recall that the weight $\tilde{\omega}(y):=\omega(x_{0}+ry)$ has the same $A_{2}$ constant than $\omega$. If we let $\tilde{c}=\omega(\{u>0\}\cap(B_{1}\setminus B_{r}(x_{0})))$ and
$\tilde{f}_{\varepsilon}(t):=r^{-n}f_{\varepsilon}(r^{n}t+\tilde{c})=\frac{1}{\varepsilon}\left(t+\frac{\tilde{c}-m}{r^{n}}\right)^{+}\qquad\hbox{for}~t\geq 0$
then, by changing variables, we get that
[…]
$=\delta^{2}r^{n-2}\bigg[\int_{B_{1}}|\nabla\tilde{u}|^{2}\tilde{\omega}\,dy+\delta^{-2}r^{2}\tilde{f}_{\varepsilon}(\tilde{\omega}(\{\tilde{u}>0\}\cap B_{1}))\bigg].$
Therefore, $\tilde{u}$ is a local minimizer in $B_{1}$ of the functional
(3.5) $\tilde{J}_{\varepsilon}(v,\tilde{\omega},B_{1})=\int_{B_{1}}|\nabla v|^{2}\tilde{\omega}\,dx+\eta\tilde{f}_{\varepsilon}(\tilde{\omega}(\{v>0\}\cap B_{1}))$
where $\eta=(r/\delta)^{2}>0$, and $\tilde{\omega}$ is an $A_{2}$ weight.
Theorem 3.4.
Let $0<\alpha<1$ be as in Theorem 2.3. There exist constants $0<\eta_{0}<1$ and $0<\lambda<1/2$ depending only on $n$, $\tilde{\omega}]_{A_{2}}$ and $\varepsilon$ such that, for any $0<\eta<\eta_{0}$ and any local minimizer $\tilde{u}$ of (3.5) satisfying
$\frac{1}{\tilde{\omega}(B_{1})}\int_{B_{1}}|\tilde{u}|^{2}\tilde{\omega}\,dx\leq 1$
we have
$|\tilde{u}(x)-\tilde{u}(0)|\leq C|x|^{\alpha/2}$
for all $x\in B_{\lambda}$, where $C>0$ depends only $n$, $\tilde{\omega}]_{A_{2}}$, $\eta_{0}$ and $\lambda$.
Proof.
We begin by setting up the constants. Let
$C_{\ast}:=\frac{C_{P,\tilde{\omega}}}{4\varepsilon}D$
where $C_{P,\tilde{\omega}}$ is the Poincaré constant in (2.3) and $D$ is the doubling constant in Lemma 2.1(2)(2) for $\tilde{\omega}$. First, we choose a radius $0<\lambda<1/4$ such that
(3.6) $4\bar{C}^{2}(D+1)\lambda^{2\alpha}\leq\frac{1}{2}\lambda^{\alpha}$
where $\bar{C}$ and $\alpha$ are the constants given in Theorem 2.3. Next, we fix $0<\eta_{0}<1$ satisfying
(3.7) $\eta_{0}\leq\min\left\{\frac{1}{C_{\ast}\lambda^{2(1-\alpha/2)}},\frac{c_{1}}{4C_{\ast}}\lambda^{\alpha}\left(\frac{|B_{\lambda}|}{|B_{1/2}|}\right)^{2}\right\}$
where $c_{1}>0$ is the constant from Lemma 2.1(1)(1) for $\tilde{\omega}$.
We prove that there exists a sequence of real numbers $\{a_{j}\}_{j\geq 0}$ such that
$(i)_{j}$ $\frac{1}{\tilde{\omega}(B_{\lambda^{j}})}\int_{B_{\lambda^{j}}}|\tilde{u}-a_{j}|^{2}\tilde{\omega}\,dx\leq\lambda^{j\alpha}$
and
$(ii)_{j}$ $|a_{j}-a_{j-1}|\leq\bar{C}\lambda^{(j-1)\frac{\alpha}{2}}$
for all $j\geq 0$, with the convention that $a_{0}=a_{-1}=0$. From here, the conclusion follows using standard arguments (see, for instance, [Stinga_2024]) as Campanato’s characterization of Hölder spaces also holds for metric measure spaces with doubling measures (see [Gorka]).
[…]
$u_{k}(x):=\frac{\tilde{u}(\lambda^{k}x)-a_{k}}{\lambda^{\frac{k\alpha}{2}}}.$
As in (3.4), $u_{k}$ is a local minimizer in $B_{1}$ of
$J_{\varepsilon}^{k}(v,\omega_{k},B_{1}):=\int_{B_{1}}|\nabla v|^{2}\omega_{k}\,dx+\eta\lambda^{2k(1-\frac{\alpha}{2})}f_{\varepsilon}^{k}(\omega_{k}(\{v+\lambda^{-k\alpha/2}a_{k}>0\}\cap B_{1}))$
where $\omega_{k}(x):=\tilde{\omega}(\lambda^{k}x)$ has the same $A_{2}$ constant as $\tilde{\omega}$ and $f_{\varepsilon}^{k}(t):=\lambda^{-kn}\tilde{f}_{\varepsilon}(\lambda^{kn}t+\tilde{c}_{k})$, with $\tilde{c}_{k}=\tilde{\omega}(\{\tilde{u}>0\}\cap(B_{1}\setminus B_{\lambda^{k}}))$.
Let $h_{k}$ be defined as $h_{k}=u_{k}$ on $B_{1}\setminus B_{1/2}$ and extended into $B_{1/2}$ as the unique weak solution to
(3.8) $\begin{cases}\mathrm{div}(\omega_{k}(x)\nabla h)=0&\hbox{in}~B_{1/2}\\ h=u_{k}&\hbox{on}~\partial B_{1/2}.\end{cases}$
By Theorem 2.3, $h_{k}$ is locally Hölder continuous in $B_{1/2}$.
We claim that
(3.9) $\frac{1}{\omega_{k}(B_{1/2})}\int_{B_{1/2}}|u_{k}-h_{k}|^{2}\omega_{k}\,dx\leq C_{\ast}\eta\lambda^{2k(1-\frac{\alpha}{2})}.$
For this, first notice that $h_{k}$ is a competitor for $u_{k}$ so
$\int_{B_{1/2}}|\nabla u_{k}|^{2}\omega_{k}\,dx\leq\int_{B_{1/2}}|\nabla h_{k}|^{2}\omega_{k}\,dx\\ +\eta\lambda^{2k(1-\frac{\alpha}{2})}\big(f_{\varepsilon}^{k}(\omega_{k}(\{h_{k}>-a_{k}\lambda^{-k\alpha/2}\}\cap B_{1})-f_{\varepsilon}^{k}(\omega_{k}(\{u_{k}>-a_{k}\lambda^{-k\alpha/2}\}\cap B_{1})\big).$
Since $f^{k}_{\varepsilon}$ is Lipschitz with Lipschitz constant $1/\varepsilon$, we get the following, where we use the doubling property of $\omega_{k}$ (see Lemma 2.1(2)(2)) in the second line:
$\int_{B_{1/2}}|\nabla u_{k}|^{2}\omega_{k}\,dx$ $\leq\int_{B_{1/2}}|\nabla h_{k}|^{2}\omega_{k}\,dx+\frac{\eta\lambda^{2k(1-\frac{\alpha}{2})}}{\varepsilon}\omega_{k}(B_{1})$
$\leq\int_{B_{1/2}}|\nabla h_{k}|^{2}\omega_{k}\,dx+\frac{\eta\lambda^{2k(1-\frac{\alpha}{2})}}{\varepsilon}D\omega_{k}(B_{1/2}).$
From here, by the Poincaré inequality (2.3) (note that $C_{P,\omega_{k}}=C_{P,\tilde{\omega}}$) and by using $u_{k}-h_{k}$ as test function in (3.8),
$\int_{B_{1/2}}|u_{k}-h_{k}|^{2}$ $\omega_{k}\,dx\leq\frac{C_{P,\omega_{k}}}{4}\int_{B_{1/2}}|\nabla(u_{k}-h_{k})|^{2}\omega_{k}\,dx$
$=\frac{C_{P,\omega_{k}}}{4}\int_{B_{1/2}}\big(|\nabla u_{k}|^{2}-\nabla u_{k}\cdot\nabla h_{k}\big)\omega_{k}\,dx$
$\leq\frac{C_{P,\omega_{k}}}{4}\int_{B_{1/2}}\big(|\nabla h_{k}|^{2}-\nabla u_{k}\nabla h_{k}\big)\omega_{k}\,dx+\frac{C_{P,\omega_{k}}}{4}\frac{\eta\lambda^{2k(1-\frac{\alpha}{2})}}{\varepsilon}D\omega_{k}(B_{1/2})$
$=C_{\ast}\eta\lambda^{2k(1-\frac{\alpha}{2})}\omega_{k}(B_{1/2})$
and (3.9) is proved. Now, since by induction hypothesis and the doubling property we have
$\int_{B_{1}}|u_{k}|^{2}\omega_{k}\,dx\leq D\omega_{k}(B_{1/2}),$
by applying (3.9), we can estimate
(3.10) $\int_{B_{1/2}}|h_{k}|^{2}\omega_{k}\,dx$ $\leq 2\int_{B_{1/2}}|h_{k}-u_{k}|^{2}\omega_{k}\,dx+2\int_{B_{1/2}}|u_{k}|^{2}\omega_{k}\,dx$
$\leq 2(C_{\ast}\eta\lambda^{2k(1-\frac{\alpha}{2})}+D)\omega_{k}(B_{1/2})$
$\leq 2(D+1)\omega_{k}(B_{1/2}),$
where we used that $\eta\leq\eta_{0}$ and (3.7).
To prove the induction step, let $a_{k+1}:=a_{k}+\lambda^{k\frac{\alpha}{2}}h_{k}(0)$. Then, by using that $\tilde{\omega}(B_{\lambda^{k+1}})=\lambda^{kn}\omega_{k}(B_{\lambda})$ and (3.9), Theorem 2.3, (3.10) and Lemma 2.1(1)(1),
$\frac{1}{\tilde{\omega}(B_{\lambda^{k+1}})}\int_{B_{\lambda^{k+1}}}$ $|u-a_{k+1}|^{2}\tilde{\omega}\,dx=\frac{\lambda^{k\alpha}}{\omega_{k}(B_{\lambda})}\int_{B_{\lambda}}|u_{k}(x)-h_{k}(0)|^{2}\omega_{k}\,dx$
$\leq\frac{2\lambda^{k\alpha}}{\omega_{k}(B_{\lambda})}\Bigg[\int_{B_{\lambda}}|u_{k}-h_{k}|^{2}\omega_{k}\,dx+\int_{B_{\lambda}}|h_{k}-h_{k}(0)|^{2}\omega_{k}\,dx\Bigg]$
$\leq\lambda^{k\alpha}\Bigg[\frac{2C_{\ast}\eta\lambda^{2k(1-\frac{\alpha}{2})}\omega_{k}(B_{1/2})}{\omega_{k}(B_{\lambda})}+4\bar{C}^{2}(D+1)\lambda^{2\alpha}\Bigg]$
$\leq\lambda^{k\alpha}\Bigg[\frac{2C_{\ast}\eta_{0}|B_{1/2}|^{2}}{c_{1}|B_{\lambda}|^{2}}+4\bar{C}^{2}(D+1)\lambda^{2\alpha}\Bigg]$
$\leq\lambda^{(k+1)\alpha}$
where in the last line we used (3.6) and (3.7). This proves $(i)_{j}$ for $j=k+1$. To see that $(ii)_{j}$ holds for $k+1$, we note that $|a_{k+1}-a_{k}|=|h_{k}(0)|\lambda^{k\alpha/2}\leq\bar{C}\lambda^{k\alpha/2}$. ∎
Theorem 3.5 (Regularity).
Any minimizer $u$ for ($P_{\varepsilon}$) is locally Hölder continuous in $B_{1}$ and, for any compact set $K\subset B_{1}$,
$u]_{C^{\alpha/2}(K)}\leq C\|u\|_{L^{2}(B_{1},\omega)}$
where $0<\alpha<1$ is as in Theorem 2.3 and $C>0$ depends only on $K$, $n$ and $\omega]_{A_{2}}$.
Proof.
Let $\eta_{0}>0$ be as in Theorem 3.4. Fix $0<\eta<\eta_{0}$ and $x_{0}\in B_{1}$. Define
$\tilde{u}(x)=\frac{u(x_{0}+rx)}{\delta}\qquad\hbox{for}~x\in B_{1}$
where
$r=\min\left(\frac{\operatorname{dist}(x_{0},\partial B_{1})}{10},\Bigg(\frac{\eta}{\omega(B_{1})}\int_{B_{1}}|u|^{2}\omega\,dx\Bigg)^{1/2}\right)\qquad\hbox{and}\qquad\delta=\frac{r}{\sqrt{\eta}}.$
Then $\tilde{u}$ satisfies the hypotheses of Theorem 3.4, so that $|\tilde{u}(y)-\tilde{u}(0)|\leq C|y|^{\alpha/2}$ whenever $|y|<\lambda$. By rescaling back to $u$, the conclusion follows. ∎
Corollary 3.6.
If $u\geq 0$ is a local minimizer of (3.1) then $\mathrm{div}(\omega(x)\nabla u)=0$ in the weak sense in the open set $\{u>0\}\cap B_{1}$.
Proof.
As a consequence of the local Hölder regularity in Theorem 3.5, $\{u>0\}$ is an open set in $B_{1}$. By Lemma 3.2, $u$ is a subsolution in $B_{1}$. To check that $u$ is a supersolution in $\{u>0\}$, consider a nonnegative $\psi\in C^{\infty}_{c}(\{u>0\})$ and any $t>0$. Since $\{u>0\}=\{u+t\psi>0\}$,
$0\leq\frac{1}{2t}(J_{\varepsilon}(u+t\psi,\omega,B_{1})-J_{\varepsilon}(u,\omega,B_{1}))=\int_{B_{1}}\omega\nabla u\cdot\nabla\psi\,dx+\frac{t}{2}\int_{B_{1}}|\nabla\psi|^{2}\omega\,dx$
and, by letting $t\to 0$,
$\int_{\{u>0\}}\omega\nabla u\cdot\nabla\psi\,dx\geq 0.\qed$
4. Behavior for small $\varepsilon$
In this section, we prove the existence part of Theorem 1.1. By the results of Section 3, the regularity properties of the minimizer stated in Theorem 1.1 hold.
The main result of this section is the following.
Theorem 4.1.
Let $\omega$ be an $A_{2}$ weight satisfying either the isoperimetric inequality (1.2) or the lower bound (1.3). Then there exists $0<\varepsilon_{*}<1$ such that if $u_{*}\in H_{g}^{1}(B_{1},\omega)$ is a nonnegative minimizer for ($P_{\varepsilon}$) as in Section 3 then
$\omega(\{u_{*}>0\}\cap B_{1})=m.$
In particular, $u_{\ast}$ is a minimizer for (1.1).
Proof.
Let $u\in H_{g}^{1}(B_{1},\omega)$ be a minimizer of ($P_{\varepsilon}$) as in Section 3. Without loss of generality, we may assume, up to rescaling, that $\inf_{\partial B_{1}}g:=\gamma\geq 1$.
Suppose that $\omega(\{u>0\}\cap B_{1})<m$. Since $u$ is locally Hölder continuous, there exists $x_{0}\in\partial\{u>0\}\cap B_{1}$ and $r>0$ sufficiently small such that
(4.1) $0<\omega(\{u>0\}\cap B_{r/2}(x_{0}))\leq\omega(B_{r}(x_{0}))<m-\omega(\{u>0\}).$
Next, let $v\in H^{1}(B_{1},\omega)$ such that $\mathrm{div}(\omega(x)\nabla v)=0$ in the weak sense in $B_{r}(x_{0})$ and $v=u$ in $B_{1}\setminus B_{r}(x_{0})$. As $\{v>0\}\cap B_{1}=(\{v>0\}\cap B_{r}(x_{0}))\cup((\{u>0\}\cap B_{1})\setminus B_{r}(x_{0}))$, we get
$\omega(\{v>0\}\cap B_{1})\leq\omega(B_{r}(x_{0}))+\omega((\{u>0\}\cap B_{1})\setminus B_{r}(x_{0}))<m.$
Hence,
$J_{\varepsilon}(v,\omega,B_{1})$ $=\int_{B_{1}}|\nabla v|^{2}\omega\,dx$
$=\int_{B_{r}(x_{0})}|\nabla v|^{2}\omega\,dx+\int_{B_{1}\setminus B_{r}(x_{0})}|\nabla u|^{2}\omega\,dx$
$<\int_{B_{1}}|\nabla u|^{2}\omega\,dx$
since $u\geq 0$ cannot be a solution to $\mathrm{div}(\omega(x)\nabla u)=0$ in $B_{r}(x_{0})$ by virtue of (4.1) and the Harnack inequality. This is a contradiction to the minimality of $u$. Thus,
$\omega(\{u>0\}\cap B_{1})\geq m.$
To prove the conclusion assume, for the sake of contradiction, that for all $\varepsilon>0$ we have $\omega(\{u>0\}\cap B_{1})>m$. Then, for all $t>0$ small enough (depending on $u$) there holds
$\omega(\{u>t\}\cap B_{1})>m.$
Define
$u_{t}=\begin{cases}\displaystyle\frac{(u-t)^{+}}{1-t}&\hbox{if}~u\leq 1\\ u&\hbox{if}~u\geq 1.\end{cases}$
Since $u_{t}\in H_{g}^{1}(B_{1},\omega)$,
$J_{\varepsilon}(u,\omega,B_{1})\leq J_{\varepsilon}(u_{t},\omega,B_{1})$
which implies
$\int_{B_{1}}\big(|\nabla u|^{2}-|\nabla u_{t}|^{2}\big)\omega\,dx\leq\frac{1}{\varepsilon}\big(\omega(\{u_{t}>0\}\cap B_{1})-\omega(\{u>0\}\cap B_{1})\big).$
Since
$\int_{B_{1}}\big(|\nabla u|^{2}-|\nabla u_{t}|^{2}\big)\omega\,dx=\int_{\{0<u<t\}}|\nabla u|^{2}\omega\,dx+\frac{t^{2}-2t}{(1-t)^{2}}\int_{\{t<u<1\}}|\nabla u|^{2}\omega\,dx$
and $\{u_{t}>0\}=\{u>t\}$, we obtain
$\int_{\{0<u<t\}}|\nabla u|^{2}\omega\,dx+\frac{1}{\varepsilon}\omega\big(\{0<u<t\}\cap B_{1}\big)\leq\frac{2t-t^{2}}{(1-t)^{2}}\int_{B_{1}}|\nabla u|^{2}\omega\,dx.$
Note that
$\int_{B_{1}}|\nabla u|^{2}\omega\,dx\leq J_{\varepsilon}(u,\omega,B_{1})\leq\int_{B_{1}}|\nabla v_{r}|^{2}\omega\,dx$
where $v_{r}$ is the function constructed in (3.2). Thus,
(4.2) $\int_{\{0<u<t\}}|\nabla u|^{2}\omega\,dx+\frac{1}{\varepsilon}\omega\big(\{0<u<t\}\cap B_{1}\big)\leq\frac{2t-t^{2}}{(1-t)^{2}}C$
where $C>0$ is a constant independent of $\varepsilon$. Next, we write the left hand side of (4.2) as
$\int_{\{0<u<t\}}|\nabla u|^{2}\omega\,dx+\frac{1}{\varepsilon}\omega\big(\{0<u<t\}\cap B_{1}\big)=\int_{\{0<u<t\}}\big(|\nabla u|^{2}+\varepsilon^{-1}\big)\omega\,dx.$
To estimate this term, recall that $u\in W^{1,1}(B_{1})$ is locally Hölder continuous by Theorem 3.5. Then we can apply the co-area formula for Sobolev functions from [maly-swanson-ziemer, Theorem 1.4] to get
$\int_{\{0<u<t\}}\big(|\nabla u|^{2}+$ $\varepsilon^{-1}\big)\omega\,dx=\lim_{N\to\infty}\int_{\{0<u<t\}\cap\{1/N<|\nabla u|<N\}}\big(|\nabla u|^{2}+\varepsilon^{-1}\big)\omega\,dx$
$=\lim_{N\to\infty}\int_{\{0<u<t\}\cap\{1/N<|\nabla u|<N\}}|\nabla u|\bigg(|\nabla u|+\frac{\varepsilon^{-1}}{|\nabla u|}\bigg)\omega\,dx$
$=\lim_{N\to\infty}\int_{0}^{\infty}\int_{\{0<u<t\}\cap\{1/N<|\nabla u|<N\}\cap\{u=s\}}\bigg(|\nabla u|+\frac{\varepsilon^{-1}}{|\nabla u|}\bigg)\omega\,d\mathcal{H}^{n-1}\,ds$
$=\int_{0}^{t}\int_{\{u=s\}}\bigg(|\nabla u|+\frac{\varepsilon^{-1}}{|\nabla u|}\bigg)\omega\,d\mathcal{H}^{n-1}\,ds.$
Consequently, and using that $x+\frac{\varepsilon^{-1}}{x}\geq 2\sqrt{\varepsilon^{-1}}$ for all $x>0$,
(4.3) $\frac{2t-t^{2}}{(1-t)^{2}}C$ $\geq\int_{0}^{t}\int_{\{u=s\}}\bigg(|\nabla u|+\frac{\varepsilon^{-1}}{|\nabla u|}\bigg)\omega\,d\mathcal{H}^{n-1}\,ds$
$\geq 2\sqrt{\varepsilon^{-1}}\int_{0}^{t}\int_{\{u=s\}}\omega\,d\mathcal{H}^{n-1}\,ds.$
Assume first that $\omega$ satisfies the isoperimetric inequality (1.2). Then, from (4.3),
$\frac{2t-t^{2}}{(1-t)^{2}}C\geq\frac{2\sqrt{\varepsilon^{-1}}}{C_{0}}\int_{0}^{t}\omega(\{u>s\}\cap B_{1})^{\frac{n-1}{n}}\,ds.$
After multiplying by $1/t$ and letting $t\to 0$, by the Lebesgue differentiation theorem, we arrive at
$CC_{0}\sqrt{\varepsilon}\geq\omega(\{u>0\}\cap B_{1})^{\frac{n-1}{n}}>m^{\frac{n-1}{n}}.$
By choosing $\varepsilon>0$ small enough, we reach a contradiction.
Assume next that $\omega$ is bounded below by $\tau$ as in (1.3). Then, by applying this assumption to (4.3) and the isoperimetric inequality from [maly-swanson-ziemer],
$\frac{2t-t^{2}}{(1-t)^{2}}C$ $\geq 2\tau\sqrt{\varepsilon^{-1}}\int_{0}^{t}\mathcal{H}^{n-1}(\{u=s\}\cap B_{1})\,ds$
$\geq 2\tau\sqrt{\varepsilon^{-1}}C_{n}\int_{0}^{t}|\{u>s\}\cap B_{1}|^{\frac{n-1}{n}}\,ds.$
After multiplying by $1/t$ and letting $t\to 0$, we get
$\frac{\sqrt{\varepsilon}}{C_{n}\tau}\geq|\{u>0\}\cap B_{1}|^{\frac{n-1}{n}}.$
Since we are assuming that $\omega(\{u>0\}\cap B_{1})>m$, by applying Lemma 2.1(3)(3), we can estimate
$|\{u>0\}\cap B_{1}|\geq\bigg(\frac{m}{C_{2}\omega(B_{1})}\bigg)^{\frac{1}{S_{D}}}|B_{1}|=:\tilde{m}$
and obtain
$\frac{\sqrt{\varepsilon}}{C_{n}\tau}\geq\tilde{m}^{\frac{n-1}{n}}$
which again gives a contradiction by choosing $\varepsilon>0$ sufficiently small.
We have shown that there exists $0<\varepsilon_{\ast}<1$ such that if $u_{\ast}\in H^{1}_{g}(B_{1},\omega)$ is a nonnegative minimizer for ($P_{\varepsilon}$) then $\omega(\{u_{\ast}>0\}\cap B_{1})=m$. We claim that $u_{\ast}$ is a minimizer for (1.1). Clearly, $u_{\ast}\in K_{0}$ and, for any $v\in K_{0}\subset H^{1}_{g}(B_{1})$,
$\int_{B_{1}}|\nabla u_{\ast}|^{2}\omega\,dx\leq J_{\varepsilon_{\ast}}(u_{\ast},\omega,B_{1})\leq J_{\varepsilon_{\ast}}(v,\omega,B_{1})=\int_{B_{1}}|\nabla v|^{2}\omega\,dx$
as claimed. ∎